Mother Zeta has two types of deposits. Deposits of both the magnitude of the first two deposits. The first deposit rate 10% p.a. and 12% second deposit. If the total interest earned from the two deposits of Rp 33.4 million a year, how much of each deposit?
1
Expert's answer
2014-10-10T11:54:02-0400
A time deposit is a money deposit at a banking institution that cannot be withdrawn for a certain "term" or period of time (unless a penalty is paid).
The formula of simple interest is:
FV = PV*(1 + i*t), I = PV*i, where PV - initial sum, i - interest rate, t - number of periods, interest earned.
In our case:
I1*0.1 + PV2*0.12 = 33.4
I1 = 334 - 10*PV2
If the initial sums of deposits are different, there is not enough information to calculate them.